OpenStudy (anonymous):

Simplify 1/4(g - 16) using the Distributive Property.

OpenStudy (experimentx):

g/4 - 16/4 = g/4 - 4

OpenStudy (experimentx):

or rather i would say 1/4*g - 1/4*16 first

Directrix (directrix):

1/4(g - 16) = (1/4) (g) - (1/4) (16) = (1/4) (g) - 4 = g/4 - 4

OpenStudy (kymber):

 :]